OpenStudy (anonymous):

Add or subtract as indicated.5x/x+1 + 6/x-1 - 10/x^2-1

OpenStudy (anonymous):

@Rowa

OpenStudy (anonymous):

to add or substract fractions we must have the same denominator on every fraction

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so you understand?

OpenStudy (anonymous):

a little my answers are A. 5x - 4 / x +1 B. 5x - 4/ x -1 C. x+1/ x - 1 D . 5x/ x - 1

OpenStudy (anonymous):

agent0smith will be able to explain it better

OpenStudy (agent0smith):

You should start learning to use the equation editor, like this: \[\Large \frac{ 5x }{ x+1 } + \frac{ 6 }{x-1 }- \frac{ 10 }{x^2 -1 }\] First step will be to factor the denominator on the far right, the x^2-1... notice it's a difference of two squares. This will help to find the common denominator for all three fractions.

OpenStudy (agent0smith):

X^2-1 factors to (x+1)(x-1): \[\Large \frac{ 5x }{ (x+1) } + \frac{ 6 }{(x-1) }- \frac{ 10 }{(x+1)(x-1) }\] Now look at the denominators on the left. What is "missing" from them, that would make the denominator the same as the (x+1)(x-1)?

OpenStudy (anonymous):

x^2 - 1 ?

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

well.. i don't understand how you got that

OpenStudy (anonymous):

but remember, we need to have the same denominator for all fractions to add or substract

OpenStudy (agent0smith):

Look at (x+1) on the right. What is missing from that, that'd make it the same as (x+1)(x-1)? compare the two: (x+1) (x+1)(x-1) what's missing from the first?

OpenStudy (anonymous):

(x - 1)

OpenStudy (agent0smith):

Good. So now we have to multiply both the numerator and denominator of it like so: \[\Large \frac{ 5x (x-1) }{ (x+1) (x-1)} + \frac{ 6 }{(x-1) }- \frac{ 10 }{(x+1)(x-1) }\] Now what's missing from the middle fractions denominator, x-1, that'll make it the same as the other two?

OpenStudy (anonymous):

(x+1)

OpenStudy (agent0smith):

Good. \[\Large \frac{ 5x (x-1) }{ (x+1) (x-1)} + \frac{ 6(x+1) }{(x+1)(x-1) }- \frac{ 10 }{(x+1)(x-1) }\] Now you can add it all together over the common denominator, (x+1)(x-1). And simplify the numerators of those two fractions, by distributing.

OpenStudy (anonymous):

i add 5x(x-1) + 6 ( x +1) - 10?

OpenStudy (agent0smith):

Yes. But you have to distribute first. So you have \[\Large \frac{ 5x (x-1) + 6(x+1) - 10 }{ (x+1) (x-1)} \]

OpenStudy (anonymous):

could you break that down as well

OpenStudy (agent0smith):

Just simplify the top. I gotta go, good luck finishing, from here it's easy! :)

OpenStudy (anonymous):

ok thank you !!

OpenStudy (anonymous):

@Rowa i got 5x-4/x+1 am i correct?

OpenStudy (anonymous):

i think 5x-4/x-1

OpenStudy (anonymous):

double check it to make sure

OpenStudy (anonymous):

ok thank you could you help me with another

OpenStudy (anonymous):

wish i could, but i have exams in a few days

OpenStudy (anonymous):

ok thanks anyways

OpenStudy (anonymous):

no problem ;)

OpenStudy (agent0smith):

5x-4/x-1 - That is correct. Make sure you use parentheses, though... (5x-4)/(x-1)